A spark is generated in an automobile spark plug when there is an electric potential of 3000 V across the electrode gap. If 75 W of power is generated in a single spark that lasts 100 ns, what is the instantaneous current across the gap?

a) 5 mA

b) 7.5 mA

c) 12 mA

d) 25 mA

e) 32 mA
Answer #1

Power=V*I, where V is voltage and I is current.

Here,V=3000 volts and power=75 W.

So,75=3000*I

=>I=75/3000=0.025 A=25 mA(option d)